OpenStudy (anonymous):

6x - 4(2x + 8) = 16 distribute then solve for x

OpenStudy (ganpat):

x = 8

mathslover (mathslover):

\[\huge{6x-4(2x+8)}\] again break down this equation into : 6x and -4(2x+8) solve for \(\huge{-4(2x+8)=-4(2x)-4(8)=-8x-32}\)

mathslover (mathslover):

now add 6x what you get ? \[\huge{-8x-32+6x=-2x-32}\] Now equate this as 16 \[\huge{-2x-32=16}\] \[\huge{-2x=16+32}\] \[\huge{-2x=48}\] \[\huge{x=-24}\]